If 3 sin 2 A + 2 sin 2 B = 1 and 3 sin 2A − 2 sin 2B = 0, where A and B are acute angles then find the value of (A +2B)


If 3 sin2 A + 2 sin2 B = 1 and 3 sin 2A − 2 sin 2B = 0, where A and B are acute angles then find the value of (A +2B)


Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear Nirabhra, 

sin2B=3/2sin2A …(i)

And from the first equation,

3sin2A=1-2sin2B=cos2B …(ii)

cos (A + 2B) = cos A. cos 2B – sin A.sin 2B

3cosA.sin2A-32sinA.sin2A

3cosA.sin2A-3sin2AcosA =0

cos (A+2B)=0

∴ A+2B=π2,3π2 …(iii)

Given that < A < π/2 and 0 < B < π/2

⇒ 0<A+2B<π+π2

⇒ 0<A+2B<3π2 …(iv)

From (iii) and (iv),

A+2B=π2

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

kaushik B
5 years ago
There is a little typo error in the solution. (square is showing as 2A)
It is given 3Sin2A+2Sin2B=1 => 3Sin2A = 1-2Sin2B => 3Sin2A = Cos 2B.....(i)
3Sin2A= 2Sin2B (given) => 3/2Sin2A=Sin2B…...(ii)
We need to find Cos(A+2B) =?   Cos(A+2B) = cosAcos2B – SinASin2B = cosA(3Sin2A)- sinA(3/2Sin2A)
= 3SinA*(1/2)*2SinAcosA – 3/2SinAsin2A = (3/2)SinAsin2A-(3/2)SinASin2A = 0= Sin(π/2) therefore
A+2B= π/2